What is the total person-time at risk over the 3 years (2000 to 2003)?

What was the cumulative incidence of HIV among study participants in January 2003?

In January 2000, 5000 men and women aged 18-30 years were recruited from a semi-urban area near Durban (South Africa) and assessed for human immunodeficiency virus (HIV). Participants who did not test positive for HIV, at baseline in 2000, were followed-up and assessed again 3 years later in January 2003 (only those who previously had tested negative for HIV were tested again at follow-up).

The following results were found:

• In 2000, at baseline: 700 participants were found to have HIV when tested

• In 2003: a further 75 cases of HIV were identified.

• These people were diagnosed at some stage during the 3 years from 2000 to 2003, the exact date of diagnosis was not available.

• Over the 3 years between 2000 and 2003: 25 people who were diagnosed with HIV at baseline in 2000 died

• By January 2003: 500 people without HIV had either died, or could not be found; no dates were known.

Key assumptions:

• HIV was acquired randomly during the 3 year period. That is, those who acquired HIV were at risk for half the time of those who did not acquire HIV or 1.5 years

• LTFU and dead were at risk for half of the 3 years or 1.5 years
Answer #1

#. The total person-time at risk over the 3 years (2000 to 2003) is calculated by multiplying the number of people participating in the study with the number of years .

Number of people participating in the study = 5000

Number of years of the study = 3 years

Total person -time at risk = 5000× 3 = 15000 times

#. Cumulative incidence of HIV among study participants in January 2003 :-

Cumulative incidence = Number of new cases of disease during a given period of time / Total population at risk

Number of new cases in 2003 = 75 new cases of HIV in a 3year period .

Total population at risk = 5000 people at risk

Cumulative incidence = 75 / 5000

= 0.015

15 cases per 1000 men and women during 3years